OpenStudy (anonymous):

From net earnings of $740 per month, Ginger Elliott must spend $200 for her portion of the rent on an apartment she shares with two friends. What percent of her net income is her rent payment? A. 32 percent B. 27 percent C. 38 percent D. 73 percent Please help me

OpenStudy (anonymous):

why are people not helping me;(

OpenStudy (help!!!!):

LOL

OpenStudy (help!!!!):

use \[\frac{ change }{ original }\times \frac{ x }{ 100 }\]

OpenStudy (help!!!!):

so what is our change?

OpenStudy (help!!!!):

@QueenZ.I.M365

myininaya (myininaya):

rent is 200 a month and she gets 740 a month so what percentage of of her income does she spend on rent it should be a low percentage because she still has 540 free to spend on whenever else which is clearly bigger than 200 so when you see percentage think that you need to put a fraction together of the info given if i said what percentage of her monthly income is not her rent I would say change/original * 100% so see if you can figure from that what her percentage on rent per month is?